OpenStudy (anonymous):

prove that 2 root 3/5 is irrational. plzzzzzzz solve this question

OpenStudy (jack1):

it cant be written as a fraction of 2 whole numbers... so it's irrational

OpenStudy (anonymous):

is it \[2\sqrt{3/5}\] ? or the 5 is outside

OpenStudy (jack1):

2 root 3/5 =sqrt 2.4 and 2.4 is rational number ( 12/5) sqrt 2.4 = 1.549193338....... and continues on perhaps infinately

OpenStudy (jack1):

there's no way you can write a non repeating decimal as a fraction of 2 numbers

OpenStudy (anonymous):

anwys assume whatever the number you've written is a rational and come up with a contradiction using that statement

OpenStudy (tkhunny):

Jack is demonstrating that it is so, but not yet proving that it is so. The usual way is by contradiction. If \(2\sqrt{\dfrac{3}{5}}\) is rational, it can be expressed as a ratio of two integers in LOWEST FORM. This last part means numerator and denominator have no common factor. So, if it's rational, we should be able to write: \(2\sqrt{\dfrac{3}{5}} = \dfrac{p}{q}\), where p and q are the reduced integers with no common factor. After a little algebra, we get: \(3\cdot q^{2} = 5\cdot 2^{2}\cdot p^2\) What can you say about this expression?